OpenStudy (anonymous):

3x+3/4< or equal to-5/11x

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

3x+3/4 <= -5/11x 3x+5/11x <= -3/4 33/33x+15/33x <= -3/4 48/33x <= -3/4 x<= (-3/4)*(33/48) x<= -.52 or you can simplify the above fraction if you want a fraction answer.

OpenStudy (anonymous):

awesome thank you i see how u go it

OpenStudy (anonymous):

yw